Solve the system of equation by the addition method
3x—3y=-12
5x+2y=-6

Answer by Boreal(15235)

3x—3y=-12
5x+2y=-6
multiply top by 2 and bottom by 3, and y will be eliminated
6x-6y=-24
15x+6y=-18
21x=-42
x=-2
substitute into first to get -6-3y=-12, -3y=-6, y=2
-10+4=-6, check in second equation
(-2, 2)